OpenStudy (vianne):

What is the value of the expression x + (–y) when x = –2.3 and y = –5.9? A. –8.2 B. –3.6 C. 3.6 D. 8.2

OpenStudy (anonymous):

Sorry, C

OpenStudy (vianne):

Could you explain how you got the answer?

OpenStudy (anonymous):

-2,3+(-(-5,9))=5,9-2,3=3,6

OpenStudy (k_lynn):

All you have to do to figure this out is substitute the values in. Since x is -2.3 and y is -5.9, you'll put -2.3 in wherever you see an x and -5.9 wherever you see a y. What will your equation look like once you've done that? @Vianne

OpenStudy (vianne):

-2.3 + (-5.9)?

OpenStudy (k_lynn):

You got it :D. So now all you have to do is add :)

OpenStudy (vianne):

Oh xD

OpenStudy (k_lynn):

:)

OpenStudy (anonymous):

c=-5,9; -c=5,9; You can write: -2,3+5,9 or 5,9-2,3

OpenStudy (vianne):

I do have another question, but I'm going to open a new one. And ok Tlvan, thanks for all of your help guys! :)

OpenStudy (k_lynn):

Sure thing.